OpenStudy (anonymous):

simplify: aa2ba3b2a2b2

OpenStudy (anonymous):

does anyone no how to do this?

OpenStudy (anonymous):

I don't. Sorry. :/ Let me search.

OpenStudy (anonymous):

Are you sure you posted the question correctly?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

"aa2ba3b2a2b2" doesn't look like a problem to me...

OpenStudy (anonymous):

the 2s and 3s are powers

OpenStudy (anonymous):

Can you rewrite it using ^. Like. a^2.

OpenStudy (anonymous):

\[a*a*2b*a*3b*2*a*2b*2 = 48 a^4 b^3 \]

OpenStudy (anonymous):

I see you have 4a \[a*a*a*a=a ^{4}\] \[2b*3b*2b=12b^{3}\] 2*2=4 4*12=48 than mytipfly you get: \[48a^{4}b ^{3}\]